OpenStudy (chycora):

Find the volume of a rectangular pyramid with a base area of 32.8 cm² and a height of 10.2 cm. Round to the nearest cubic centimeter. A. 446 cm³ B. 335 cm³ C. 1,051 cm³ D. 112 cm³

OpenStudy (chycora):

@welshfella @pandaluvs

OpenStudy (chycora):

@Baby_Bear69

OpenStudy (welshfella):

Volume of pyramid = 1/3 * height * area of the base

OpenStudy (welshfella):

= 1/3 * 10.2 * 32.8

OpenStudy (chycora):

ok it's 111.520

OpenStudy (chycora):

@welshfella

OpenStudy (welshfella):

Yes thats correct

OpenStudy (chycora):

so is d the answer

OpenStudy (anonymous):

@chycora yes, d. is the answer because the formula being used is correct: V=1/3 of Bh, where B is the area of the base and h is the height. If you round 111.520 to the nearest cubic centimeter, it is 112, so yes it is D.

OpenStudy (chycora):

k thanx